Write the equation 6x-3y=9 in slope intercept form
Answer by ejazali_syed(7)

slope intercept form
y = mx + b
so we have to convert our given equation into the above form.
given equation 6x - 3y = 9
6x - 9 = 3y ==> y = (6x - 9)/3

==> y = 6x/3 - 9/3
==> y = 2x - 3